# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

### v2

@@ 4.2 Quantitative Comparisons
User Study We conducted a user study comparing our method with other baseline methods. We randomly selected 20 sets of stylized views of 3D scenes from both the LLFF and T&T datasets, processed by various methods, and invited 25 participants. Participants were then asked to vote on the video based on four evaluation indicators: visual preference, style match level, content preservation level, and 3D consistency quality. More details are included in the supplementary document. We collected a total of 1000 votes for all evaluation indicators and presented the results in Fig. 11 in the form of boxplots. Our method, StylizedGS, outperformed other methods across all evaluation indicators, indicating visually preferable and consistent results. A p-value of 0.00001 demonstrates the statistical significance of the obtained results.
4.3 Controllable Stylization Results
Our method enables users to control stylization in three ways, including color, scale, and spatial region. The corresponding controllable stylization results are shown in Figs. 3.2, 4 and 5, respectively. Fig. 12 shows an example where a sequential series of controls are applied to a scene across different conceptual factors. In addition to the direct stylization result, we sequentially apply different controls from left to right: preserving the color of the original scene, increasing the scale of the pattern style, and adopting spatial control to apply two styles. Such flexibility allows for controllable refinement of key parameters or stylistic attributes, satisfying specific aesthetic preferences or artistic requirements and fostering enhanced artistic expression.
@@ 4.3 Controllable Stylization Results
TABLE II: Quantitative ablation study results. We report ArtFID and SSIM for our method and other ablations over 50 randomly chosen stylized cases.
@@
4.4 Ablation Study
@@
Additionally, we investigate how the proposed 3DGS filter affects the reconstructed scenes. Normally, floaters become more noticeable after the style transfer process. As shown in Fig. 14, the texture of stones and leaves extend into the air around the objects without the 3DGS filter. In contrast, the results of our method exhibit enhanced clarity, with fewer visible colored floaters. Tab. II shows the quantitative results of the ablation studies, demonstrating the contributions of all components of our method.
5 Discussion and Conclusion

### v3

@@ IV-B Quantitative Comparisons
TABLE IV: Quantitative ablation study results. We report ArtFID and SSIM for our method and other ablations over 50 randomly chosen stylized cases.
@@
User Study We conducted a comprehensive user study with 52 participants evaluating 40 stylized 3D scenes across multiple datasets. Participants compared our method against baselines using four metrics: visual preference, style fidelity, content preservation, and 3D consistency. Each participant assessed randomized method pairs alongside style references and original scenes, yielding 1000 total votes. Statistical analysis confirms our approach’s superiority across all metrics ($p\leq 0.0001$), with results demonstrating significant preference for our method in visual quality and consistency (Fig. 15). Study design ensured demographic diversity, unbiased evaluation conditions, and rigorous statistical validation, as detailed in the supplement.
IV-C Controllable Stylization Results
Our method enables users to control stylization in three ways, including color, scale, and spatial region. The corresponding controllable stylization results are shown in Figs. 3, 4 and 5, respectively. Fig. 16 shows an example where a sequential series of controls are applied to a scene across different conceptual factors. In addition to the direct stylization result, we sequentially apply different controls from left to right: preserving the color of the original scene, increasing the scale of the pattern style, and adopting spatial control to apply two styles. Such flexibility allows for controllable refinement of key parameters or stylistic attributes, satisfying specific aesthetic preferences or artistic requirements and fostering enhanced artistic expression.
IV-D Ablation Study
@@
Additionally, we investigate how the proposed 3DGS filter affects the reconstructed scenes. Normally, floaters become more noticeable after the style transfer process. As shown in Fig. 18, the texture of stones and leaves extend into the air around the objects without the 3DGS filter. In contrast, the results of our method exhibit enhanced clarity, with fewer visible colored floaters. Tab. IV shows the quantitative results of the ablation studies, demonstrating the contributions of all components of our method.
V Discussion and Conclusion
@@
Supplementary Material for StylizedGS: Controllable Stylization for 3D Gaussian Splatting
I Overview
This supplementary document accompanies our main paper and provides implementation details along with additional stylization results. It is divided into four sections: implementation details, more stylization results, additional validation experiments, and a discussion about our future work.
  • Section II provides the implementation details, including training settings, more details about the color transfer, spatial control, and user study.
  • Section III provides additional stylization results, including more controllable stylization results.
  • Section IV provides additional validation experiments, including more comparisons with baseline methods on LLFF and T&T datasets, and an additional ablation study.
  • Section V provides a discussion of our future work.
II Implementation Details
II-A Training Settings
We implement StylizedGS based on the 3DGS representation [17]. In the color transfer step, we fix the density component of the initial 3DGS and optimize it using $\mathcal{L}{rec}$ for 200 iterations. We empirically set the filtering ratio k% of the opacity to 5% and that of the scale to 8% for all scenes. We run the filter once every 100 iterations. During the stylization process, we fine-tune all parameters of $G{\theta}^{sty}$ over 800 iterations, while disabling the adaptive density control of 3DGS. We adopt a pretrained VGG-16 network [30] to extract the feature maps $F_{render}$, $F_{style}$, and $F_{content}$ used in the NNFM loss. The network consists of 5 convolution blocks, and we use the ’conv2’ and ’conv3’ blocks as the feature extractors, following the observations of the ablation study in Sec. IV.
The loss weights $(\lambda_{rec},\lambda_{sty},\lambda_{con},\lambda_{sca},\lambda_{opa},\lambda_{tv})$ are respectively set to $(0.2,2,0.005,0.05,0.05,0.02)$. The depth preservation loss weight $\lambda_{dep}$ is set to $0.01$ for all forward-facing captures and $0.05$ for all 360 captures. At each stylization training iteration, we render an image from a sampled view different from all the training views to compute losses. The Adam optimizer [78] is adopted with a learning rate exponentially decayed from $0.1$ to $0.01$.
II-B Color Transfer
In the color transferring procedure, we employ histogram matching to adjust the color distribution of the style image to align with that of the content images. We select this transformation to ensure that the mean $\mu$ and covariance $\Sigma$ of the RGB values in the recolored training views $p_{c}^{re}$ match all pixels in the style image $p_{s}$. The transformation solution $\mathbf{A}$ and $b$ are determined to satisfy the constraints:
$b$ $=\mu_{p_{c}}-\mathbf{A}\mu_{p_{s}}$ (12)
$\mathbf{A}\Sigma_{p_{s}}\mathbf{A}^{T}$ $=\Sigma_{p_{c}}$
A family of solutions for $\mathbf{A}$ that satisfies these constraints exists. We consider utilizing the principal component analysis (PCA) method. First, let the eigenvalue decomposition of a covariance matrix be $\Sigma=\mathbf{U}\Lambda\mathbf{U}^{T}$. Then, we define the matrix square-root as: $\Sigma^{1/2}=\mathbf{U}\Lambda^{1/2}\mathbf{U}^{T}$. Subsequently, the transformation is given by:
$\mathbf{A}=\Sigma_{p_{c}}^{1/2}\Sigma_{p_{s}}^{-1/2}$ (13)
The transformation solutions are also applied to the color parameter $c$ to recolor the 3DGS representation $G_{\theta}^{rec}$.
II-C Spatial Control
During our spatial control process, users can acquire region masks via point-based interaction or language-based interaction. In the language-based approach, the text prompt enables the generation of masks for different perspectives using LangSAM [57]. In the point-based approach, users assign several points on a single-view image, and our mask-tracking strategy facilitates the attainment of consistent multi-view masks.
Our mask-tracking strategy is based on the assumption that the area of the mask regions should be similar across different perspectives, which yields better performance in forward-facing scenes. Upon user specification of points $P_{0}$, SAM [56] is employed to obtain a mask $M_{0}$ for the initial view. When acquiring the mask for subsequent views, the initially specified points remain fixed. If the area of the acquired mask falls within a certain error range $\Delta s$ relative to the first mask $M_{0}$, it is assigned as the final mask for that view. If the discrepancy is significant, a search is conducted within the neighborhood pixels of the specified point to find a mask with an area close to the first mask. This process is repeated for all perspectives, with each subsequent view’s mask search beginning from the final specified point of the previous view. Algorithm 1 illustrates the general process for our mask-tracking strategy.
Input : User specified points $P_{0}$, traing view images $\{I_{i}\}_{i\in[n]}$
$\triangleright$ Initialization
[…]
$\triangleright$ Dynamically Tracking
for $i\leftarrow 1$ to n1n-1 do
[…]
      end for
    $P_{i-1}\leftarrow P$
   end while
end for
Output : $\{M_{i}\}_{i\in[n]}$
ALGORITHM 1 StylizedGS mask tracking.
II-D User Study
In the user study, we randomly selected 40 sets of stylized views of 3D scenes from both the LLFF and T&T datasets, processed by various methods. We invited 52 participants (28 males and 24 females, aged 18 to 50). Our questionnaire was carefully designed to include participants from diverse backgrounds and age groups, reducing the risk of echo chamber effects. Participants were given ample time for unbiased evaluation, ensuring a thorough and fair assessment. Initially, we presented the participants with a style image, the original scene, and two stylized videos generated by our method and a randomly chosen comparison method. Participants were then asked to vote on the videos based on four evaluation indicators: visual preference, style match level, content preservation level, and 3D consistency quality. The specific descriptions of these four indicators are as follows:
  • Visual preference: Which one do you prefer in terms of visual quality?
  • Style match level: Which one has a higher level of matching with the style image?
  • Content preservation level: Which one maintains the content of the original scene in a better way?
  • 3D consistency: Which one has better 3D consistency of the scene after stylization? 3D consistency refers to whether the scene remains consistent when the camera perspective changes in the rendered video. Generally speaking, some flickering can be observed in the scene when it is inconsistent.
These descriptions are also shown to participants at the beginning of the questionnaire.
Figure 1: Additional color control stylization results. Our approach facilitates versatile color management in stylized outputs, allowing users to retain the scene’s original hues or apply distinct color schemes from alternative style images. Users can choose to transfer the entire style, only the pattern style, or a mix of arbitrary patterns and color styles.
Figure 2: Additional scale control stylization results. Our method enables users to flexibly control the scale of basic style elements, such as adjusting the density of colorful blocks, as demonstrated in the first row.
Figure 3: Additional spatial control stylization results. By specifying masks with text prompt in the style image and content images, users can transfer different styles to desired regions with semantic correspondence. Note that the style images are realistic images, so the results are photo-realistic style transfer.
III Additional Results
Stylization. We provide additional stylization results on the T&T dataset in Fig. 4 These illustrations show our method’s ability to faithfully capture both color and pattern styles across different views.
Controllable Stylization. An advantage of our approach is the ability to perform controls over stylization. We present additional color control results in Fig. 1, scale control results in Fig. 2, and spatial control results in Fig. 3. These results demonstrate the capability of our controllable stylization method. It’s noteworthy that in spatial control, the style images used are realistic, resulting in photo-realistic style transfer. Moreover, to enhance the semantic correspondence between the style image and content images, we employ LangSAM to extract masks through text input for both style and content images, facilitating the transfer of specified styles.
IV Additional Validations
More Comparisons. We present additional qualitative comparisons with baselines on the T&T dataset in Fig. 7 and on the LLFF dataset in Fig. 5. It is evident from these comparisons that our method surpasses other techniques in effectively transferring both geometric patterns and color styles.
Figure 4: More stylization results on T&T dataset. It can be seen that our method faithfully captures both the color styles and pattern styles across different views.
Figure 5: Additional comparisons with baselines on LLFF dataset. Compared to other methods, our method excels in learning intricate and accurate geometric patterns while effectively preserving the semantic content of the original scene.
Figure 6: Ablation study of feature blocks. We visualize the stylization results of extracting features from different blocks of the VGG-16 network. It can be seen that the combination of features from ‘conv2’ and ‘conv3’ blocks is the best choice, which has better stylization results.
Figure 7: Additional comparisons with baselines on T&T dataset. It can be seen that our method has better results which faithfully capture both the color styles and pattern styles across different views.
Additional Ablation Study. In order to select appropriate feature maps in the calculation of the NNFM loss, we conduct an ablation study on the feature maps extracted from different blocks of the VGG-16 network, as shown in Fig. 6. From the comparison, it can be seen that selecting both ‘conv2’ and ‘conv3’ blocks for feature extraction preserves style details better than other alternatives.
V Future Work
Reconstructed 3DGS representation may include some floaters, which do not have large impact for scene rendering. However, they can significantly disrupt our stylization process. To address this issue, we propose a filter-based refinement approach. Nevertheless, it is worth noting, as mentioned in the limitations, that this process cannot entirely eliminate the impact of floaters.
In future work, we intend to explore the integration of recent advancements in Gaussian reconstruction methods [77, 79]. These methods incorporate regularization terms or compress the 3D volume into a series of 2D oriented planar Gaussian disks, thereby mitigating the occurrence of floaters and improving overall reconstruction quality. On the other hand, we will explore feed-forward-based 3DGS stylization methods, especially controllable stylization, which need to overcome the challenge between sparse user interactions and parameter changes of dense 3D Gaussians.
Finally, current language-based editing and stylization methods provide users with a new way of interaction. We will further explore how to implement language-based controllable stylization on 3DGS.